# A classification of Fourier summation formulas and crystalline measures

### v1

@@ 7. Proof of Theorem 1
$\mu$ $=-\sum_{\gamma}{\boldsymbol{\delta}}_{\gamma}$
$\nu_{1}$ $=2\cosh(\pi t)\mathrm{d}t$
$\nu_{2}$ $=-\frac{1}{2\pi}{\rm Re}\,\psi(\tfrac{1}{4}+i\tfrac{t}{2})\mathrm{d}t,$
then $(\mu,a)\in\rm ACM$. The difference of any two of such pairs above for two characters $\chi_{1}$ and $\chi_{2}$ real, even and primitive, but different moduli $N$, produces a pair in $\rm CM$ (this was pointed out first by Guinand [18]). Such measures are also connected with new Fourier interpolation formulae produced in [11], and these produce, under GRH, $\rm CM$-pairs $(\mu,a)$ with $\operatorname{\mathrm{supp}}(\mu)=\{\gamma\in\mathbb{R}:L(\chi,1/2+i\gamma)=0\}\cup\{\gamma_{0}\}$ and $\operatorname{\mathrm{supp}}(a)=\{\operatorname{\mathrm{sgn}}(n)\frac{\log|n|}{4\pi}\}_{n\geq 1}$, where $\gamma_{0}$ is any real such that $L(\chi,1/2+i\gamma_{0})\neq 0$.

### v3

@@ 8.5. Almost Crystalline Measures
$\nu$ $=-2\cosh(\pi t)\mathrm{d}t$
$\mu$ $=-\sum_{\gamma}{\boldsymbol{\delta}}_{\gamma}+\frac{1}{2\pi}{\rm Re}\,\psi(\tfrac{1}{4}+i\tfrac{t}{2})\mathrm{d}t$
then $(\mu,a)\in\rm ACM$. The difference of any two of such pairs above for two characters $\chi_{1}$ and $\chi_{2}$ real, even and primitive, but different moduli $N$, produces a pair in $\rm CM$, this was pointed out first by Guinand [22], and we explain how this construction fits our framework in the end of Remark 6. Such measures are also connected with the remarkable new Fourier interpolation formulae produced in [15]. Assuming all non-trivial zeros of $L(\chi,s)$ have real part $1/2$ and are simple, these new formulae produce $\rm CM$-pairs $(\mu,a)$ with $\operatorname{\mathrm{supp}}(\mu)=\{\gamma\in\mathbb{R}:L(\chi,1/2+i\gamma)=0\}\cup\{\gamma_{0}\}$ and $\operatorname{\mathrm{supp}}(a)=\{\operatorname{\mathrm{sgn}}(n)\frac{\log|n|}{4\pi}\}_{n\geq 1}$, where $\gamma_{0}$ is any real such that $L(\chi,1/2+i\gamma_{0})\neq 0$.
Acknowledgments
We thank Oleksiy Klurman and Bryce Kerr for all the conversations we had around this topic during their visit to Rio. Theorems do come out of the smoke. We thank Friedrich Littmann for the fruitful conversations and comments which inspired Theorem 5. We also thank Michael Baake and Peter Sarnak for further discussions, references and remarks after the first version of this paper. The author acknowledges support from the following funding agencies: The Office of Naval Research GRANT14201749 (award number N629092412126), The Serrapilheira Institute (Serra-2211-41824), FAPERJ (E-26/200.209/2023) and CNPq (309910/2023-4).